OpenStudy (anonymous):

what is this?? sqrt27u^3v^10w^6

OpenStudy (anonymous):

@mathlover2014 @iambatman @undeadknight26

OpenStudy (owlcoffee):

\[\sqrt{27u^3 v^10 w^6}\] Like this?

OpenStudy (anonymous):

almost. its v to the tenth

OpenStudy (owlcoffee):

okay, what do you want to with that?

OpenStudy (anonymous):

it says simplify the expression

OpenStudy (anonymous):

and assume all variables are positive

OpenStudy (owlcoffee):

\[\sqrt{27u^3v ^{10}w^6}\] Here you can take advantage of a property of radicals: \[\sqrt{u.v}=\sqrt{u}\sqrt{v}\] Si, here, we can take the square root of each factor: \[\sqrt{27u^3v ^{10}w^6}=\sqrt{27}\sqrt{u^3}\sqrt{v ^{10}}\sqrt{w^6}\]

OpenStudy (anonymous):

oh okay. yeah I remember learning that. but is that the answer?? arent there supposed to to be like, exponents and fractions and stuff?

OpenStudy (anonymous):

dont you do like (27u^3v^10x^6)^1/2

OpenStudy (owlcoffee):

Yes, but I will not give you the answer, I am not allowed to do that.

OpenStudy (anonymous):

i dont know where to go from there

OpenStudy (anonymous):

i know you cant give the answer. but would you be willing to help me? I dont know where to go after (27u^v^10w^6)^1/2

OpenStudy (owlcoffee):

If you want to take the method if expressing the main radical as a fractionary exponent, you can take advantage of a exponential property: \[(27u^3v ^{10}w^6)^{\frac{ 1 }{ 2 }}\] \[(a^n)^m=a ^{mn}\] This expands to each factor, so: \[(27^{\frac{ 1 }{ 2 }})(u ^{3\frac{ 1 }{ 2 }})(v ^{10\frac{ 1 }{ 2 }})(w ^{6\frac{ 1 }{ 2 }})\]

OpenStudy (anonymous):

and then what? @owlcoffee

OpenStudy (owlcoffee):

simplify the exponents.